In a local area the government intends to reduce the coyote population by 10% each year. If the coyote population is now estimated to be 3,000, how many coyotes will remain after 3 years?
Answer by Mathtut(3670)

3000-3000(.1)=2700
:
2700-2700(.10=2430
:
2430-2430(.10)=2187